1. An investor is pondering on whether to invest in a siomai franchise or in a printing press company. He had his staff analyze the cost benefits that he would get from the two business ventures and they came up with this report.


A. The siomai franchise would need Php 800 000 as capital. It has a 60-percent chance of profiting Php 400 000; a 15-percent chance of profiting Php 600 000; and 25-percent chance of losing the capital entirely.


B. The paper manufacturing company would need Php 800 000 as capital. It has a 35-percent chance of earning Php 1 000 000 and a 65-percent chance of losing the capital entirely.


C. Which between the two business ventures should the investor invest his Php 800 000?


1
Expert's answer
2021-05-12T07:27:23-0400

A. The siomai franchise expected profit is:"EP = 400,000*0.6 + 600,000*0.15 - 800,000*0.25 = 130,000."

B. The paper manufacturing company expected profit is:

"EP = 1,000,000*0.35 - 800,000*0.65 = -170,000."

So, the investor should invest his Php 800,000 into the siomai franchise.